2. For each of the following, compute the Present Value. Show your work.

Present Value

Years

Interest Rate

Future Value

13

10%

15,500

4

8%

51,555

29

24%

886,072

40

35%

550,165
Answer #1

1.

Present value = Future value / (1+i)n

= 15,500/ (1+0.10)13

= 15,500/(1.10)13

= 15,500/3.452271

= $4,489.80

2.

Present value = Future value / (1+i)n

= 51,555/ (1+0.08)4

= 51,555/(1.08)4

= 51,555/1.36048896

= $37,894.46

3.

Present value = Future value / (1+i)n

= 886,072/ (1+0.24)29

= 886,072/(1.24)29

= 886,072/511.951559

= $1,730.77

4.

Present value = Future value / (1+i)n

= 550,165/ (1+0.35)40

= 550,165(1.35)40

= 550,165/163,437.135

= $3.36

Present Value

Years

Interest Rate

Future Value

$4,489.80

13

10%

15,500

$37,894.46

4

8%

51,555

$1,730.77

29

24%

886,072

$3.36

40

35%

550,165

Kindly comment if you need further assistance.

Thanks‼!